OpenStudy (anonymous):

let U=<-1,-3> and V=<5,5> . Find -3u+5v

OpenStudy (anonymous):

Scalar multiplication of a vector goes like this:\[C(a,b)=(Ca,Cb)\] and addition of vectors goes like this:\[(a,b)+(c,d)=(a+c,b+d)\]What's your best guess at the value of \(-3 \vec u + 5\vec v\)?

OpenStudy (anonymous):

(22, 34)

OpenStudy (anonymous):

-3(-1,-3)+5(5,5)= (3,9)+(25,25)=3+25,9+25)=(28,34)

OpenStudy (anonymous):

So, the answer is (28,34)

OpenStudy (mathmate):

Correct! Well done!